OpenStudy (anonymous):

Ten students are applying for 3 positions on a team. the students include 4 boys (adam, alex, anthony, and arnold) and 6 girls (abbey, aurora, agnes, alice, amanda, and anna). all the students have an equal chance of being selected. find the probability that the students selected will include: a) at most 1 girl b) adam, anthony, and alice c) agnes and 2 other students

OpenStudy (anonymous):

ur in college?

OpenStudy (anonymous):

naww not in college lol :P highschool why'd u think im in college?

OpenStudy (m):

the first one should be (4/10)^3 + (6/10)^1 x (4/10)^2

OpenStudy (m):

do you have answers to confirm?

OpenStudy (anonymous):

i was working on his question until you answered it. i was simply wondering why theis person thought i was in college. he asked me a question, i was answering.

OpenStudy (m):

second one should be (6/10)^1 x (4/10)^2

OpenStudy (m):

is it with replacement or without? lol

OpenStudy (anonymous):

What do you think?

OpenStudy (m):

i guess without :P

OpenStudy (m):

yeah i always get confused with statistics

OpenStudy (m):

you may have to use hypergeometric forumula

OpenStudy (m):

wwhat i wrote is binomial

OpenStudy (anonymous):

Well, you certainly don't 'have to' - if one really wanted to they could solve this by just drawing a (massive) tree diagram, but I think we can be slightly more clever about it,

OpenStudy (m):

hmm maybe (6/10)(4/9)(3/8) x 3 + (4/10)(3/9)(2/8)?

OpenStudy (m):

crap i'm so confused with this i should just stop lol. gl with the problem.

OpenStudy (anonymous):

I don't have answers to go with this.

OpenStudy (anonymous):

anyone figure this out?

OpenStudy (anonymous):

Don't worry, you can just draw a tree-diagram with 720 branches and be sure of the answer.

OpenStudy (m):

[(6/10)(4/9)(3/8) + (4/10)(6/9)(3/8) + (4/10)(3/9)(6/8)] + (4/10)(3/9)(2/8)

OpenStudy (anonymous):

Proof?

OpenStudy (m):

wait the ordering doesn't matter so you have to divide the first part by 3? or 6?

OpenStudy (anonymous):

I preferred it in the old form you posted: (6/10)(4/9)(3/8) x 3 + (4/10)(3/9)(2/8)

OpenStudy (m):

only problem with that is you don't know girl is selected before boys

OpenStudy (anonymous):

You don't. But your answer there accounts for it by multiplying by three.

OpenStudy (m):

i don't know stat well enough to figure this out without final answer unfortunatley

OpenStudy (anonymous):

OK. If there are three boys chosen, the chance is (clearly... I hope) (4/10)(3/9)(2/8) The only other option is 1 girl and 2 boys. This can either be GBB, BGB, BBG, but the chance of each is still (6x4x3)/(10x8x9). So your answer \[\frac{(6 \times 4 \times 3) \times 3 + 4 \times 3 \times 2}{720} \] is right.

OpenStudy (anonymous):

for which part?

OpenStudy (m):

even though probability of getting boy or girl is different?

OpenStudy (m):

oh wait you already accounted for that lol

OpenStudy (anonymous):

Yes; there's still a total number of 10*9*8 = 720 in the denominator; the order of the multiplication above is irrelevant. @ azntiger627 - FACEPALM - if you honestly have no idea which part this is then you have clearly done nothing on this work, and do not deserve help until at least attempt it.

OpenStudy (anonymous):

I just came back from dinner lol

OpenStudy (anonymous):

Wasting time eating... when you have Maths to do!? That's a cardinal sin in my book.

OpenStudy (anonymous):

lol sorry, now I'm rdy, so which part were u guys talking about?

OpenStudy (anonymous):

I'll tell you. But you have to guess first.

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

ok so what did u find out so far?

OpenStudy (anonymous):

Nothing :(

OpenStudy (m):

Newton already solved part (a) for you above. but i think you should go over permutations and combinations before trying to solve the question above. good luck!

OpenStudy (anonymous):

You solved it first :(

OpenStudy (anonymous):

D) ?

OpenStudy (anonymous):

for d, i'm guessing (4C2)(6C1) / 10C3

OpenStudy (anonymous):

Very funny, Mr; there is no D.

OpenStudy (anonymous):

lol b I meant

OpenStudy (anonymous):

is that correct though?

OpenStudy (anonymous):

That equal 3/10 - does that seem correct?

OpenStudy (anonymous):

yea I got 3/10, is it correct though?

OpenStudy (anonymous):

There are 720 possibilities, and you think 216 contain the same three people? (they don't)

OpenStudy (anonymous):

what do u think then?

OpenStudy (anonymous):

I don't 'think' anything; I know what the answer is.

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

i will look at it soon, im tied up atm

OpenStudy (anonymous):

There names all begin with A, which is annoying, so just called them ABC. There are 6 ways of arranging them: ABC ACB etc. And 10 x 9 x 8 possible ways in total. Continue from there.

OpenStudy (anonymous):

Their names* UGGGGGGGGGH grammar fail.

OpenStudy (anonymous):

6/720?

OpenStudy (anonymous):

¬_¬ maybe